### v1

# ♪  Something Just Like TRuST ♪ thanks: An allusion to Coldplay’s “Something Just Like This.” : Toxicity Recognition of Span and Target

Abstract
Toxicity in online content, including content generated by language models, has become a critical concern due to its potential for negative psychological and social impact. This paper introduces TRuST, a comprehensive dataset designed to improve toxicity detection that merges existing datasets, and has labels for toxicity, target social group, and toxic spans. It includes a diverse range of target groups such as ethnicity, gender, religion, disability, and politics, with both human/machine-annotated and human/machine-generated data. We benchmark state-of-the-art large language models (LLMs) on toxicity detection, target group identification, and toxic span extraction. We find that fine-tuned models consistently outperform zero-shot and few-shot prompting, though performance remains low for certain social groups. Further, reasoning capabilities do not significantly improve performance, indicating that LLMs have weak social reasoning skills.
♪  Something Just Like TRuST ♪ thanks: An allusion to Coldplay’s “Something Just Like This.” : Toxicity Recognition of Span and Target
Disclaimer: Due to the topic studied here, the paper contains offensive words.
Berk Atil  Namrata Sureddy  Rebecca J. Passonneau Penn State University {bka5352,nqs5685,rjp49}@psu.edu
1 Introduction
Offensive and toxic content is pervasive in social media and online forums. Because large language models (LLMs) are trained on online data, they learn to generate toxic output Gehman et al. (2020); Hartvigsen et al. (2022). Exposure to toxic content might lead to lack of empathy, prejudice Pluta et al. (2023), stress, and depression Saha et al. (2019). Social groups that might need special attention such as people with developmental disorders can be more vulnerable to toxic content Kim et al. (2023). Furthermore, LLMs are being used in high-stakes conversational applications such as education Yu et al. (2025); de Araujo et al. (2025) and healthcare Yeo et al. (2024, 2025). Exposure to offensive language can lead to issues with self-esteem, anger, and anxiety Kansok-Dusche et al. (2023). Within the NLP community there has been increasing interest in toxicity detection Davidson et al. (2017); Rawat et al. (2024); Duan et al. (2025); Khurana et al. (2025) and toxicity mitigation Suau et al. (2024); Ermis et al. (2024); Pozzobon et al. (2023a); Li et al. (2024); Liu et al. (2021). However, existing toxicity datasets differ in format, leading to lack of uniformity in benchmarking. We build on existing datasets to create TRuST to improve benchmarking, and to provide a more comprehensive picture of online use of toxic language.
Figure 1: Histogram of distinct tokens within toxic spans (x-axis) ordered by their total count (y-axis). Note that the figure has a vertical "tear" at around 150 on the x-axis to show continuation out beyond 500 distinct tokens, but with counts too low to be visible. The most common 25 words are shown descending on a diagonal.
Existing toxicity datasets for English differ in size, source, and annotation scheme. They can range from about 5K examples Zampieri et al. (2023) to four times that size Davidson et al. (2017); Mathew et al. (2021) for human-generated posts from social media (typically twitter), or an order of magnitude larger for AI-generated datasets Hartvigsen et al. (2022). Davidson et al. (2017) introduced a toxicity scheme for hate speech, offensive language, or neither that has been widely adopted Almohaimeed et al. (2023); Mathew et al. (2021), although binary labels for offensiveness have also been used and combined with hierarchical schemes for the target of offense Zampieri et al. (2019). Categorizations of target include higher level classes for individual, group or other Zampieri et al. (2019), as well as very fine-grained labels, thus the THOS dataset Almohaimeed et al. (2023) has 31 targets (called topics). Some datasets also include identification of the spans of words that support identification of toxicity or offensiveness within a post or comment Zampieri et al. (2023); Mathew et al. (2021); Pavlopoulos et al. (2022). We review existing datasets and their annotation schemes to arrive at a synthesis that includes toxicity, target and span. In particular, we define a closed set of social group targets that includes half a dozen higher level classes and over 20 fine-grained subgroups. Our efforts produce TRuST, a large scale dataset of over 300K examples, created by merging many existing datasets that represent diverse social media sources as well as AI-generated examples. We also train human annotators to label a subset of 11,500 examples, reaching interannotator agreement scores in line with previous work.
To investigate the benefits of TRuST, we carry out three activities: an analysis of the characteristics of toxic language in social media (cf. Fig. 1); benchmark tests of LLMs on toxicity detection; automatic annotation of all the TRuST examples apart from the manually annotated subset, using the best method from our benchmarking tests. We benchmark LLMs using zero-shot, few-shot, and Chain-of-Thought prompts and compare LLMs to pre-trained language models (PLM) that we fine-tune. We find that fine-tuned PLMs perform somewhat better than LLMs. We also find that performance remains low on certain social groups, suggesting that social reasoning in LLMs needs improvement.
Our contributions are: (1) TRuST, a comprehensive dataset with toxicity, target social group, and toxic span labels; (2) a comprehensive definition of toxicity based on a synthesis of previous work; (3) benchmarking of SOTA LLMs, and their comparison with previous SOTA models.
2 Related Work
2.1 Definitions of Toxicity
Since the perception of toxicity depends on factors such as context, background, demographics, it is subjective. This has led to work that provides parallel annotations of toxicity from different demographic groups Mostafazadeh Davani et al. (2024). Further, types of toxicity annotation have ranged from narrower categories such as hate speech to broader types of offense.
Davidson et al. (2017) define hate speech HS as hatred towards a social group or humiliation/insult to the members of a group. Additionally, HS might include threats or reference to violence. Nockleby (2000) defines HS as any communication that disparages a person or a group on the basis of a social characteristic such as race, color, ethnicity, gender, sexual orientation, nationality, religion, or other characteristic. These works generally agree that HS is directed at a target.
Other types of toxic content go beyond hate speech. Dorn et al. (2024) state that toxic content includes using slurs in a pejorative way, attacking or criticizing a minority, promoting hate speech, negative stereotypes, etc. The definition used in PerspectiveAPI Lees et al. (2022) for toxic content is “rude, disrespectful or unreasonable language that is likely to make someone leave a discussion”. Kumar et al. (2021) propose that toxicity involves identity-based attacks such as racism, bullying, threats of violence, sexual harassment, and includes emotional harm and hate speech.
One issue in the literature about toxicity is that some terms are used interchangeably, or one of them is seen as a subcategory of the other. Therefore, there is a lack of consensus which leads to confusion and inconsistency Khurana et al. (2022). Moreover, some works consider every content with slurs or swear words as toxic, which may not be the case (e.g. you are f***ing beautiful).
We take toxicity to comprise three categories: hate speech, abusive language, and sexual harassment. Hate speech is defined as offensive and discriminatory discourse towards a group or an individual based on characteristics such as race or religion, thus always has a target. It includes negative stereotyping (negative traits and characteristics attributed to a group), racism (discrimination against or negative attitudes towards individuals/groups belonging to a certain race), sexist language (discrimination or behaviors that foster stereotypes based on a gender), and discrimination based on sexual orientation. Abusive language is content with inappropriate words such as profanity or disrespectful terms for people based on sociodemographic characteristics. It includes psychological threats (expressions of an intent of harms such as humiliation, causing distress, constant criticism). Our last category is sexual harrasment which includes unwelcome sexual moves, requests of sexual favors, or other unwanted physical/verbal behaviors of a sexual nature towards someone. In our work, toxic language often has a target, but can also involve use of offensive words in an aggressive fashion without targeting a specific social group, e.g., “honestly? I can handle kpop stans dragging armys but just stay the f**k away from bts they’ve done lit rally nothing to y’all.” Our annotation instructions (see Appendix C) include a binary label for toxicity that covers these three categories.
2.2 Datasets
Davidson et al. (2017) produced one of the earliest datasets focusing on hate speech and offensiveness in social media. Their dataset has around 25k tweets with three-way labels for hate speech, offensive language or neither, assigned by crowdworkers. OLID Zampieri et al. (2019) has 14k tweets and hierarchical offensiveness labels. The first level is offensive or not. Offensive examples are then labeled as targeting a group or not. The last level is for the targeted offensive tweets for whether the target is an individual, social group, or any other category such as an organization.
Inclusion of annotations for target social group supports deeper understanding of which are the groups that are targeted most often, and other bias-related issues. TBO Zampieri et al. (2023) is a recent dataset that has target and toxic span annotations, both of which support explainability. However, their target annotation involves identifying specific words in toxic language samples, rather than more general categories, which limits their utility. The THOS hate speech dataset Almohaimeed et al. (2023) has two-level categories for target: the first is more high level such as country or religion (6 classes), and the second is more detailed belonging to the first level categories (31 classes) such as a term for a country (China), practitioner of a religion (Muslim), or racial group (Black). Similar to THOS, we employ two levels with 8 higher levels (including No Target), 5 of which are broken down into two or more subclasses.
Another annotation type that supports explainability is the span of words that constituted the offense, as in ToxicSpans Pavlopoulos et al. (2022). HateXplain Mathew et al. (2021) has target group annotations based on religion, gender etc. and rationales that are the spans on which annotators’ labeling decision is based. The rational annotation can contain words for target group. Our set of targets is more diverse in the number of target groups.
The cited works mostly focus on explicit hate speech; implicit hate speech is harder to detect. ElSherief et al. (2021) propose a new dataset with 19k tweets with an implicit versions of hate and toxicity (e.g., irony). Social Bias Frames Sap et al. (2020) creates a formalism that models pragmatic frames. They have annotations for targeted group, implied statement, offensiveness, use of in-group language etc. Alongside datasets with social media texts, some works create data using generative language models, such as ToxiGen Hartvigsen et al. (2022). The latter focuses on implicit hate speech; they generate both toxic and benign examples for 13 minority groups using GPT-3 Brown et al. (2020). Their dataset has 274k examples and labels are based on the prompts used for GPT-3 with the assumption that if the prompt is toxic, the generated text is also toxic. FairPrism Fleisig et al. (2023b) is another machine-generated dataset that builds on ToxiGen and Social Bias Frames. They collect human annotations on stereotyping and demeaning harms, target social group, and providing non-aggregated annotations to support fairness.
Some argue that the context, author, and reader are important factors for perception of toxic language Cowan and Hodge (1996); Nieto and Boyer (2006). This motivates Zhou et al. (2023) to use GPT3.5 to build CobraFrames, with contextual annotations for speaker and listener, target group, and emotional reactions or power dynamics between the speaker and listener.
2.3 Prediction of Toxicity, Targets and Spans
The SOTA for toxicity detection is PerspectiveAPI Lees et al. (2022), in which multilingual BERT Devlin et al. (2019) is first trained, then multiple single-language convolutional neural networks are distilled. However, it has been criticized for having only moderate correlation with humans Welbl et al. (2021); Schick et al. (2021), and over-reliance on surface patterns such as swear words, resulting in a high rate of false positives Rosenblatt et al. (2022). Another issue is that they sometimes release updates without notification or explanation of improvements, making previous findings questionable Pozzobon et al. (2023b).
A common approach for target social group is a multi-layer perceptron classifier whose input is embeddings of examples from a pretrained encoder, such as BERT Mathew et al. (2021). LLMs have also been used Zhou et al. (2023).
Target Count (%) Toxic % T. Count (%) T. Toxic % No target 4121 ($35.96$) $38.26$ 358 (36.46) 37.99 Ethnicity 2050 ($17.78$) $55.10$ 170 (17.31) 51.76 black 723 ($6.24$) $\mathbf{74.90}$ 64 (6.52) 70.31 white 278 ($2.45$) $46.43$ 21 (2.14) 38.10 asian 272 ($2.34$) $47.80$ 23 (2.34) 43.48 native 169 ($1.51$) $\mathbf{32.63}$ 16 (1.63) $\mathbf{18.75}$ chinese 157 ($1.32$) $37.95$ 8 (0.81) 37.50 o. ethnicity 129 ($1.13$) $43.66$ 11 (1.12) $\mathbf{72.73}$ mexican 114 ($0.97$) $38.52$ 9 (0.92) 44.44 arab 105 ($0.90$) $65.49$ 7 (0.71) 57.14 latino 103 ($0.93$) $45.30$ 11 (1.12) 27.27 Politics 1281 ($11.05$) $63.12$ 103 (10.49) 72.82 Gender 1152 ($9.92$) $49.56$ 87 (8.86) 55.17 lgbtq+ 521 ($4.50$) $50.00$ 38 (3.87) 55.26 woman 492 ($4.24$) $\mathbf{50.75}$ 38 (3.87) $\mathbf{57.89}$ man 121 ($1.02$) $48.44$ 9 (0.92) $\mathbf{44.44}$ o. gender 18 ($0.17$) $\mathbf{14.29}$ 2 (0.20) 50.00 Religion 1112 ($9.77$) $58.62$ 99 (10.08) 53.53 muslim 528 ($4.58$) $55.11$ 41 (4.18) $\mathbf{41.46}$ jewish 474 ($4.21$) $\mathbf{66.60}$ 49 (4.99) $\mathbf{65.31}$ o. religion 110 ($0.98$) $\mathbf{40.65}$ 9 (0.92) 44.44 Other 825 ($7.22$) $51.49$ 78 (7.94) 52.56 other 466 ($4.07$) $\mathbf{54.97}$ 44 (4.48) $\mathbf{56.82}$ refugee 188 ($1.66$) $\mathbf{41.15}$ 17 (1.73) $\mathbf{41.18}$ middle east 171 ($1.49$) $53.48$ 17 (1.73) 52.94 Country 545 ($4.73$) $29.60$ 50 (5.09) 26.00 o. country 357 ($3.11$) $\mathbf{30.10}$ 34 (3.46) $\mathbf{32.35}$ US 188 ($1.61$) $\mathbf{28.57}$ 16 (1.63) $\mathbf{12.50}$ Disability 412 ($3.57$) $30.22$ 37 (3.77) 29.73 Total 11498 $47.89$ 982 47.35
Table 1: Statistics for our human annotated data showing the total count (and percentage of the total) for each higher level or lower-level social group, and the percentage of each that are labeled toxic. Lower-level groups with the highest and lowest proportion of toxic texts are in red and green font, respectively. The last two columns are for the test set (T.). In targets, o. means other, native means native american.
Toxic span prediction is usually treated as a multi-token classification problem where a label is predicted for each token He et al. (2024). SpanBERT Joshi et al. (2020), a pretrained model for high quality embeddings for predicting spans of text, is the SOTA encoder for this task. While there is little work on using LLMs for toxic span extraction, it has been done for Romanian toxic span identification using GPT4 Paraschiv et al. (2023).
3 Dataset
TRuST re-annotates data from ToxicSpan (CC-BY 4.0) Pavlopoulos et al. (2022), HateXplain (CC-BY 4.0)Mathew et al. (2021), TBO (CC-BY 4.0) Zampieri et al. (2023), Thos (CC-BY 4.0) Almohaimeed et al. (2023), and ToxiGen (CC-BY 4.0) Hartvigsen et al. (2022). We observed some noise in the annotations, which is inevitable due to the subjective nature of toxicity. The ToxiGen toxicity labels, however, were particularly noisy, due to the strong assumption of assigning the prompt labels to the generated text. Where necessary, we added annotations for target group and span. TRuST therefore has three types of labels, toxicity, target group and span, throughout. The full dataset has nearly 300K examples (see Table A).
To provide a high-quality benchmark for testing toxicity models, we collected human annotations for binary toxicity, target social group, and toxic spans for a subset of over 10k examples. We have 24 target social groups, including “no target”. Note that we separate “Chinese” from Asian because it includes a significantly large number of examples. After testing different models on prediction of the three types of labels, we annotate the rest of the data with the best-performing models.
3.1 Human Annotation Procedure
To collect the manual annotations, we hired six undergraduate students studying computer science or data science who have some experience with data analysis; they were paid \$$\$10$/hour. We recruited students from different backgrounds (Indian, Chinese, White) and demographics to have more diversity, given the cultural and demographic differences in toxicity Mostafazadeh Davani et al. (2024); Fleisig et al. (2023a). They were provided with detailed instructions (see Appendix C). To ensure good quality, we did three iterations of annotator training. We should note that by the time we collected toxic span annotation, which was done last, only three of the original six annotators were still available. Our inter-annotator agreement scores using Krippendorf’s $\alpha$ Krippendorff (2013) are 0.56, 0.66, and 0.55 for toxicity, target social group and spans, respectively. Agreement scores in previous work (Krippendorff’s $\alpha$ unless otherwise noted) are, for toxicity: 0.46 Mathew et al. (2021), 0.51 Sap et al. (2020) and 0.64 Hartvigsen et al. (2022); for target social group: 0.50 Sap et al. (2020); for spans: 0.55 Cohen’s kappa Pavlopoulos et al. (2022)). To calculate agreement for spans, we incorporate MASI Passonneau (2006), a distance metric for inter-annotator agreement on sets, based on a weighted Jaccard, where the weight is higher for set subsumption than for set intersection, which in turn is higher than for disjunction.
After annotators were trained, they first annotated target social group for every example (including no target), followed by toxicity. During the toxicity annotation, they also double-checked the previously labeled target social group. The examples where a previous target social group annotation was considered incorrect were labeled by all students, where we applied majority voting to get the final label. Span annotation came last, during which students did a sanity check on the toxicity label, where those judged incorrect were again relabeled by the three remaining annotators, with majority voting for the final label. Notably, span annotation identified either individual words that led to the judgment that the example was toxic, or applied to the entire sentence.
3.2 Toxicity and Target Group Labels
Table 1 shows the total count for each higher- or lower-level target group in the human-annotated subset, along with the proportion of these that are labeled toxic. Most target groups are ethnicity-based; the least frequent target group is for "disability." In general, we have almost a balanced dataset for toxic versus non-toxic content (47.37% of the examples are toxic), but the rate varies greatly within each group; for example, 75% of the examples for the social group "black" are toxic, while only 33% are for "native american."
3.3 Span Labels
The total number of examples with annotation of the toxic span is 5,506 (47.89% of 11,498, per Table 1), where some examples have multiple spans. To calculate descriptive statistics on span tokens, we first apply stemming. We merge the strings “ni**a” and “ni**er”. The mean length of spans is 1.91 words, the median is 1, and the maximum is 11. In 33% of cases with span annotation, the span constituted the entire sentence. There are 7065 spans in total, among which 1334 are unique. The histogram in Figure 1 shows the span tokens occurring more than once is highly skewed. We show the 25 most common span tokens, with “ni**er”, “f**k”, “stupid”, “idiot”, and “b**ch” at the top. Some words are specific to particular groups such as “black” or “kike” but others such as “kill” or “stupid” are used more generally.
Model Accuracy Precision Recall F1 PerspectiveAPI 0.50 0.48 0.77 0.59 RoBERTa 0.79 0.76 0.82 0.79 BERT 0.79 0.77 0.79 0.78 GPT4o 0.75 0.70 0.85 0.77 Sonnet 0.77 0.72 0.83 0.77 Llama70b 0.77 0.71 0.86 0.78 Llama8b 0.73 0.66 0.88 0.76 Reasoning and CoT D. Llama70b 0.75 0.69 0.84 0.76 D. Llama8b 0.70 0.63 0.90 0.74 o4-mini 0.78 0.71 0.90 0.79 GPT4o-cot 0.74 0.73 0.71 0.72 Llama8b-cot 0.73 0.67 0.84 0.75 Llama70b-cot 0.75 0.72 0.75 0.74 Sonnet-cot 0.73 0.73 0.67 0.70
Table 2: Toxicity detection results. D. models are R1 distilled versions.
4 Experiments
Our experiments utilize the human-annotated data of 11,498 examples divided into validation, training and test as follows. The validation set is 495 examples, with randum selection plus enforcing a minimum of 5 examples per target group. The randomly selected test set has 982 examples. We compare performance of multiple baselines on the three tasks of detection of toxicity, target social group, and span detection for toxic examples. We first compare PLMs versus zero-shot LLMs on each task in turn, including a custom model in the case of toxicity detection, followed by a subsection on whether prior knowledge of target social group improves toxicity detection. Then we compare these results to the use of reasoning models and in-context learning. We report accuracy, precision, recall and F1 for the first two tasks, but omit accuracy for toxic span detection, as the exact matching needed for accuracy is less informative than the breakdown of F1 into recall and precision on toxic span words. The final subsection presents an error-analysis.
PLM baselines for toxicity and target social group use BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) with linear classifier layers, and for span prediction we use SpanBERT Joshi et al. (2020) (see appendix B). We include four LLMs: GPT4o OpenAI et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and Llama3.1 (70b and 8b) Grattafiori et al. (2024) (see Appendix E for the prompt). We use temperature=0, set the seed, and use the default values for others for determinism, although it is not guaranteed Atil et al. (2025). For toxicity, we also include PerspectiveAPI Lees et al. (2022), a neural network that provides a probability that an example is toxic. We used the validation set to identify the best probability for the binary class cutoff, which was 0.20.
4.1 Toxicity
The toxicity results in the top of Table 2 show that fine-tuned PLMs perform slightly better than the LLMs. Except for Llama8b, LLMs perform similarly. Surprisingly, PerspectiveAPI’s accuracy is random. All models have higher recall which is preferable here, where false negatives are worse than false positives.
Model Accuracy Precision Recall F1 BERT 0.76 0.68 0.82 0.72 RoBERTa 0.73 0.63 0.81 0.70 GPT4o 0.75 0.67 0.78 0.70 Sonnet 0.74 0.62 0.75 0.67 Llama70b 0.65 0.55 0.68 0.58 Llama8b 0.48 0.15 0.15 0.15 Reasoning and CoT o4-mini 0.72 0.58 0.63 0.59 D. Llama70b 0.69 0.24 0.27 0.25 D. Llama8b 0.61 0.17 0.17 0.17 GPT4o-cot 0.75 0.68 0.75 0.70 Llama8b-cot 0.40 0.07 0.06 0.06 Llama70b-cot 0.68 0.50 0.58 0.53 Sonnet-cot 0.75 0.62 0.70 0.65
Table 3: Target group prediction results.
Model Accuracy Precision Recall F1 BERT 0.80 0.77 0.82 0.79 RoBERTa 0.78 0.74 0.81 0.77 GPT4o 0.66 0.64 0.65 0.64 Sonnet 0.71 0.65 0.70 0.66 Llama70b 0.57 0.46 0.48 0.44 Llama8b 0.49 0.07 0.07 0.07 Reasoning and CoT o4-mini 0.67 0.45 0.47 0.45 D. Llama70b 0.64 0.17 0.18 0.17 D. Llama8b 0.56 0.07 0.07 0.07 GPT4o-cot 0.68 0.56 0.61 0.58 Llama8b-cot 0.44 0.05 0.05 0.05 Llama70b-cot 0.57 0.31 0.31 0.31 Sonnet-cot 0.70 0.55 0.60 0.57
Table 4: Higher level target results.
4.2 Target Social Group
Similar to Zampieri et al. (2023), for two baselines we train a neural network with a linear classifier layer on top of BERT or RoBERTa encoder. As above, we also test the same SOTA LLMs (cf. appendices F and G for the prompts).
Similar to toxicity detection results, fine-tuning a model slightly outperforms LLMs, as seen for the fine-grained target results in Table 3. However, Llama models do not perform as well as GPT4o and Sonnet. The F1 for Llama8b is especially low, showing an inability to understand social targets of toxicity. Table 4 shows the results for the higher-level target groups. Interestingly, LLMs perform worse at this task than the fine-grained target groups (9% decrease for GPT4o, 3% decrease for Sonnet etc.). Confusion matrices for GPT4o and Sonnet (the two top-performing LLMs), show that GPT4o mixes “other” and “ethnicity” predictions with “no target”, and Sonnet mixes “no target” with “ethnicity”. Additionally, both mix “ethnicity” with “other”. This indicates “other” as a high level target is more confusing than the fine-grained targets.
Model Precision Recall F1 SpanBERT 0.72 0.71 0.70 GPT4o 0.55 0.79 0.65 Sonnet 0.66 0.45 0.53 Llama70b 0.66 0.22 0.33 Lama8b 0.48 0.48 0.48 Reasoning and Cot o4-mini 0.63 0.40 0.49 D. Llama70b 0.45 0.87 0.59 D. Llama8b 0.43 0.58 0.4 Sonnet-cot 0.63 0.5 0.56 GPT4o-cot 0.55 0.63 0.59 Llama70b-cot 0.68 0.46 0.54 Llama8b-cot 0.68 0.16 0.26
Table 5: Toxic span prediction results.
(a) Toxicity few-shot results.
(b) Target group few-shot results.
(c) Higher Target Group Few-Shot Results.
(d) Toxic Span Few-Shot Results.
Figure 2: Few-Shot Comparison Figures
4.3 Toxic Span
Table 5 shows that SpanBERT outperforms the LLMs. GPT4o’s F1 approaches SpanBERT’s, but the other LLMs do much worse.
4.4 Target Group and Toxicity Detection
We conduct experiments to assess whether prior knowledge about target social group could improve toxicity detection. For the RoBERTa/BERT-based models, we either add this information at the text level as “The target social group is <social group>” or at the embedding level. For LLMs, we use two types of prompting. In the first, we assign the social group persona to the model; combining persona with self-correction has been effective Xu et al. (2024). In the second, we include the social group in the text of the prompt. We find no differences between the methods (cf. Table 7 of the Appendix).
4.5 Reasoning Models
We also experiment with chain-of-thought (CoT) Wei et al. (2023) and the reasoning models o4-mini and R1 Distilled Llama70b/8b . Although reasoning helps in science or logic Jaech et al. (2024); DeepSeek-AI (2025); Zhang et al. (2024); Wei et al. (2023), it does not improve detection of toxicity or target group, and gives mixed results on span prediction. CoT helps Llama70b and Sonnet, but not Llama8b and Gpt4o. Reasoning usually increases recall, which indicates that models predict more tokens less precisely.
Figure 3: Toxicity detection accuracy by target group
Figure 4: F1 scores of the models on the span data labeled as “all sentence” (x-axis) vs others (specific spans are found, y-axis). Reasoning models, few-shot prompted models, and zero-shot models are labeled with a different color.
4.6 In-Context Learning
Few-shot learning improves LLM performance in many tasks, such as question answering Brown et al. (2020). There are many strategies for selecting few-shot examples. Previous work suggested that similarity to the example in question is effective Paraschiv et al. (2023); Liu et al. (2022); Zebaze et al. (2024) was borne out in our early experiments. We use Linq-Embed-Mistral Kim et al. (2024) to create embeddings, and retrieve similar examples from our training data. Figure 2 shows results for few-shot up to six, for 4 non-reasoning models. For toxicity detection, examples help only for GPT4o, and requires at least three examples to surpass zero-shot. For toxic span prediction, one example is enough to increase the performance for Llama models, whereas Sonnet requires at least two examples. For target group prediction, few-shot learning helps all models, and the improvements are more dramatic, e.g., from 0.48 to 0.64 for Llama8b with one example. However, apart from prediction of target group, LLMs still do not surpass the performance of fine-tuned PLMs.
4.7 Error Analysis
The best results reported above (accuracy; F1) are from the RoBERTa classifier for toxicity (0.79; 0.79), the BERT classifier for fine-grained target (0.76; 0.72) and course-grained target (0.80; 0.79); SpanBERT for toxic spans (F1 of 0.70). In general, LLMs do better at toxicity detection than they do on the other two tasks, and reasoning/CoT helps very little. Here we look into differences across social groups for toxicity and target group prediction, whether span detection is easier when the whole sentence is the toxic span, and performance differences on machine-generated text.
Differences across social groups. For the error analysis, we broke down the accuracy results into accuracy per target social group for toxicity and target group detection. The Figure 3 accuracy heatmap for toxicity detection shows wide variation across target groups, ranging from as low as 41% up to 100%. Apart from the categories for "other gender" and "other ethnicity," where most models have very high performance, all models show non-uniform performance across groups. In some cases, pairs of models correlate in their performance (e.g. Spearman correlations are 0.89 for GPT4o and Sonnet; 0.87 for o4-mini and Llama70b; 0.83 for Sonnet and Llama70b), but in no case model performance correlates well with data support.
We also found large differences in target group accuracy across target social groups (see Appendix Figure 6). For example, Llama8b has a 42% difference in accuracy for "black" versus "white." The models struggle in categories such as "other country" or "other gender," possibly because these categories are necessarily more heterogenous.
Span detection for sentences versus subsentence spans. As mentioned in Section 3.1, toxic spans sometimes consist of the entire sentence (35% of the test set). Figure 4 plots F1 for full sentence toxic spans on the x-axis by subsentence spans on the y-axis. Most models, including SpanBERT, have much higher performance on full-sentence toxic spans. Distilled Llama models are far worse at detecting toxic subsentence spans.
LLM-generated text. Our test data composition is 40% GPT-3 generated (from ToxiGen). Breaking down accuracy by human versus LLM origin shows 20% greater accuracy on the GPT-3 generated text.
5 Conclusion
To further research into identifying and mitigating toxicity in LLMs, we have presented TRuST, a dataset of over 298K examples, where about 38% are toxic, and have been labeled for 24 fine-grained target social groups (including "no target"), and toxic spans. A subset of over 11K examples were manually annotated by a diverse group of annotators whose inter-annotator agreements on the three labeling tasks were higher than reported in previous work. We benchmarked 13 methods for detection accuracy and F1 on the three aspects of toxicity. Unsurprisingly, the PLM classifiers outperformed the LLMs. Neither CoT nor reasoning models led to higher LLM performance, thus current methods to improve LLM reasoning do not address social reasoning well, if at all.
The 287K examples that were not human-annotated have been automatically annotated using the best performing models: a RoBERTa PLM for toxicity, a BERT PLM for target group, and SpanBERT for toxic span. This dataset will be released if the paper is accepted for publication. We offer TRuST as a contribution toward more robust methods for toxicity detection, and perhaps to support development of corpora and methods to mitigate toxicity, such as unlearning Chen and Yang (2023); Liu et al. (2024), which to our knowledge has not yet been applied to toxicity.
6 Limitations
The work presented here carries out only a preliminary investigation of baseline methods for automatic identification of toxicity, target social group and toxic span detection. The LLMs methods did not explore sophisticated prompt engineering. Although the size of the dataset is competitive, it is not sufficiently large to have separate annotations for some important subgroups. Although we attempted to recruit a pool of annotators that was socially diverse, this was limited due to lack of funds to recruit more than six annotators.
Appendix A Statistics on the Whole Dataset
Target H. Count (%) H. Toxic % M. Count (%) M. Toxic % C. Count (%) C. Toxic %
No Target 3618 ($36.10$) $38.92$ 73211 ($25.40$) $24.03$ 76829 ($25.76$) $24.73$
Ethnicity 1763 ($17.59$) $56.21$ 68962 ($23.93$) $48.02$ 70725 ($23.71$) $48.22$
[…]
white 244 ($2.43$) $49.18$ 8691 ($3.02$) $51.29$ 8935 ($3.00$) $51.24$
asian 238 ($2.37$) $49.16$ 12961 ($4.50$) $45.36$ 13199 ($4.43$) $45.43$
native american 143 ($1.43$) $36.36$ 8817 ($3.06$) $34.04$ 8960 ($3.00$) $34.07$
chinese 139 ($1.39$) $37.41$ 10242 ($3.55$) $41.53$ 10381 ($3.48$) $41.47$
other ethnicity 108 ($1.08$) $38.89$ 5156 ($1.79$) $35.47$ 5264 ($1.76$) $35.54$
mexican 95 ($0.95$) $35.79$ 6221 ($2.16$) $42.65$ 6316 ($2.12$) $42.54$
arab 88 ($0.88$) $65.91$ 2267 ($0.79$) $70.27$ 2355 ($0.79$) $70.11$
latino 82 ($0.82$) $52.44$ 5044 ($1.75$) $49.15$ 5126 ($1.72$) $49.20$
Politics 1132 ($11.30$) $62.54$ 9624 ($3.34$) $57.41$ 10756 ($3.61$) $57.95$
Gender 1000 ($9.98$) $50.20$ 34280 ($11.89$) $42.67$ 35280 ($11.83$) $42.88$
[…]
woman 431 ($4.30$) $50.58$ 14839 ($5.15$) $47.94$ 15270 ($5.12$) $48.02$
man 102 ($1.02$) $49.02$ 5248 ($1.82$) $43.45$ 5350 ($1.79$) $43.55$
other gender 11 ($0.11$) $9.09$ 1531 ($0.53$) $14.30$ 1542 ($0.52$) $14.27$
Religion 966 ($9.64$) $60.97$ 30381 ($10.54$) $50.65$ 31347 ($10.51$) $50.97$
[…]
jewish 407 ($4.06$) $68.06$ 14765 ($5.12$) $58.23$ 15172 ($5.09$) $58.49$
other religion 91 ($0.91$) $42.86$ 5127 ($1.78$) $31.44$ 5218 ($1.75$) $31.64$
Other 716 ($7.14$) $51.96$ 14771 ($5.12$) $43.73$ 15487 ($5.19$) $44.11$
[…]
refugee 160 ($1.60$) $42.50$ 4710 ($1.63$) $39.53$ 4870 ($1.63$) $39.63$
middle east 144 ($1.44$) $52.78$ 7408 ($2.57$) $42.54$ 7552 ($2.53$) $42.73$
Country 470 ($4.69$) $31.49$ 23189 ($8.05$) $25.02$ 23659 ($7.93$) $25.15$
[…]
united states 162 ($1.62$) $33.33$ 9080 ($3.15$) $25.55$ 9242 ($3.10$) $25.69$
Disability 357 ($3.56$) $30.53$ 24332 ($8.44$) $27.05$ 24689 ($8.28$) $27.11$
Total 10222 48.16 288233 37.79 298255 38.14
Table 6: The statistics on the training data including machine annotated data. The second and third columns are human annotated training data (H), the fourth and fifth are machine annotated training data (M), and the last two are the combination of both (C). The results for the higher level categories for human annotation are computed by combining the fine-grained category results. For the machine generated, however, we carry out independent tests of the higher and lower categories. As a result, we find a discrepancy of about 10% of cases.
Appendix B Experimental Details
We use AdamW optimizer Kingma and Ba (2017) to fine-tune PLMs. We use a batch size of 8. We tried 0.01, 0.0001, 0.00001, 0.05,0.005,0.0005, 0.00005 for learning rate and we chose the best one for each task, we chose the best one based on the validation performance. The best learning rate is 1e-05 for all tasks and embedders. We ran the experiments on a single NVIDIA RTX A6000 GPU and it took 2 hours for each experiment.
Appendix C Annotation Instructions
Goal
Creating annotated data for toxic language, where, very generally, a sentence is toxic if it has negative stereotyping, hate speech, racism, psychological threat, sexual harassment, abusive language, sexism, discrimination based on sexual orientation, or any other type of explicit language that might hurt or affect a member of some sociodemographic group badly.
Task
Annotate these 3 categories:
  • Toxicity (binary)
  • Social target group (24 categories belonging to 6 higher level groups including other, and no target)
  • Toxic Span (words that make the sentence toxic, only for toxic sentences.)
Assumptions
  • There is no speaker information (who produced the language)
  • There is no audience information (who the speaker was addressing, not necessarily the target).
  • The toxicity label should take into account the target social group (note that the same words might be toxic for one social group target and not for another)
  • There is no context, so texts should be interpreted as they are, based on easily available readings.
  • If you see a politician’s name/party name, choose politics.
  • If the target is individual, we are not interested in so they should be labeled as no target.
  • If there are multiple targets, try to choose the dominating one. In case they are equally targeted, choose other.
Target Groups
Higher Level Target Groups We have 5 distinct higher level social targets, plus labels for “Other” category (there is a target but it doesn’t fall into these categories) and absence of a target (“No Target”).
  • Ethnicity/Nationality
  • Gender
  • Religion
  • Disability
  • Country
  • Other
  • No Target
Full Set of Target Groups We have more specific targets within higher level categories. For each category, except individual and disability, we also have other. You should use the label “other $<higher\_level\_target>$” in the case that the group belongs to $<higher\_level\_target>$ and not among the ones listed below (e.g. for the statement “Turkish people are useless”, the target is ethnicity, but there is no Turkish category so it should be other ethnicity)
  • Black, white, Asian, Mexican, Native American, Chinese, Latino, and Arab belong to ethnicity.
  • LGBTQ+, Woman, and Man belong to gender.
  • Muslim and Jewish belong to religion.
  • Disability does not have any sub-targets.
  • United States is the only country belonging to country. You should label as “other country” if it is another country.
  • Middle East and refugee belong to other.
  • Immigrants belongs to refugee
Note: Do not forget about the “other” categories within each category, except individual and disability!
Toxicity Label Criteria
For assigning positive/negative toxicity; toxic if any of the following categories, where the first five (blue font) are varieties of hate speech, the next two are examples of abusive language other than sexual harrassment, the final category.
  • Negative stereotyping: Traits, characteristics and generalizations that are negatively attributed towards a group. e.g. Turkish people are selfish
  • Hate speech: Offensive and discriminatory discourse towards a group or an individual based on characteristics such as race or religion. E.g. All Jews should be killed.
  • Racism: Usually a bias that includes acceptance of negative stereotypes, and racial discrimination (discrimination; different treatment of different age, gender, racial etc., race in this case) against individuals/groups belonging to a certain race. E.g Arabs should not be offered any jobs.
  • Sexist language: Discrimination or behaviors that foster stereotypes based on a gender. E.g. Men cannot be successful in teaching
  • Discriminating based on sexual orientation: E.g. Being lesbian should be banned.
  • Psychological threat: It is an expression of an intent of harm, such as humiliation, causing stress, or constant criticism. E.g. You always suck
  • Abusive language: Piece of text that contains inappropriate words such as swear profanity, and disrespectful nicknames. E.g It is no wonder everyone says you’re a jerk
  • Sexual harassment: Unwelcome sexual moves, sexual favor requests or other physical/verbal behaviors of a sexual nature towards someone. E.g. Your breasts are amazing.
Note: there are other situations of language use that can be perceived as toxic that do not fall under our guidelines given our assumptions, e.g., some microaggressions might be considered toxic but to identify them would require other information such as the audience, more of the overall context in which the language occurs, etc.
Toxic Spans
You should choose the word(s) that causes the toxicity of the statement (e.g. We need to ignore the needs of disabled people).
General Instructions
  • We should think as being in the target group.
  • You should consider multiple possible interpretations since almost every statement/utterance has multiple “readings”; If there is an easily available understanding that leads to being non-toxic, it should be labeled as non-toxic.
  • For the higher level target, you need to check if the opinion is towards some group, individual or no one. There might be more than one group, in that case try to choose the dominant group. Also, if it is for an individual and seems like they belong to a social group, do not try to infer a social group for them, instead label them as “individual”
Appendix D Potential Risks
Due to the nature of our work, TRuST has some potential risks of increasing toxic content. However, all text data were already public before our work. We add human and machine annotations to these texts, which should not increase the current potential risk. We hope this work will be used for good, not to make the system more harmful.
Appendix E LLM Prompt for Toxicity
System Prompt:
You are an expert in linguistic and sociological analysis. Your task is to analyze a text and determine if it is toxic or not. Toxicity here refers to negative stereotyping, hate speech, racism, sexist language, discrimination based on sexual orientation, physical or psychological threat, and abusive language. Note that there are other situations of language use that can be perceived as toxic that do not fall under our guidelines, given our assumptions For each text, at the end, respond in the following format: Toxicity Label: [toxic/non-toxic]
User Prompt:
Text: <text>
Appendix F LLM Prompt for Target Group
System Prompt:
You are an expert in linguistic and sociological analysis. Your task is to determine which social group is targeted by the sentiment or content of each text. For each text, use the provided list of social groups and always select the group that is most directly addressed or implied. If an individual not involved in politics is mentioned, or if no group is targeted, respond with ’Targeted Group: no target’. If multiple groups are targeted with no clear dominance, or if a group is certainly targeted but not clearly identified, respond with ’Targeted Group: other.’ Here are the social groups:
- no target - other country - united states (notes: must be related to the United States as a whole) - black - asian (note: ethnicity, excludes Chinese) - white - arab - latino (note: ethnicity, includes Hispanic and excludes Mexican) - native american - chinese - mexican - other ethnicity - jewish - muslim - other religion - politics (note: must reference a political figure or party) - woman - man - lgbtq+ - other gender - disability - middle east - refugee (note: includes immigrants) - other For each text, at the end, respond in the following format: Targeted Group: [Group Name] User Prompt: Text: <text>
Appendix G LLM Prompt for Higher Level Target Group
System Prompt:
You are an expert in linguistic and sociological analysis. Your task is to determine which social group is targeted by the sentiment or content of each text. For each text, use the provided high level social groups and always select the group that is most directly addressed or implied. If an individual not involved in politics is mentioned, or if no group is targeted, respond with ’Targeted Group: no target’. If multiple groups are targeted with no clear dominance, or if a group is certainly targeted but not clearly identified, respond with ’Targeted Group: other.’ Here are the social groups: - no target - country - ethnicity - religion - politics - gender (note: includes lgbtq+) - disability - other (note: includes politics that references a political figure or party) For each text, at the end, respond in the following format: Targeted Group: [Group Name]
User Prompt:
Text: <text>
Appendix H LLM Prompt for Toxic Span
System Prompt:
You are an expert in linguistic and sociological analysis. You are an expert in linguistic analysis. The task is to label toxic sequences in user-generated comments. You place the toxic sequences between [[ and ]] and do not change the text otherwise. If the overall text is toxic but not some specific words based on the meaning, you should output "all sentence". A word can be toxic because of its usage in the sentence and it may not be toxic on its own. You need review the replies before output. Be concise and exact. Answer only with JSON statements such as: "output": "lol crybaby [[faggot]]"
User Prompt:
Text: <text>
Appendix I Effect of Target Group on Toxicity Detection
Model Acc. Pre. Recall F1
RoBERTa 0.79 0.76 0.82 0.79
RoBERTa w target 0.78 0.74 0.83 0.78
GPT4o 0.75 0.70 0.85 0.77
GPT4o persona 0.75 0.73 0.74 0.74
GPT4o w target 0.73 0.74 0.68 0.71
Sonnet 0.77 0.72 0.83 0.77
Sonnet persona 0.72 0.72 0.68 0.70
Sonnet w target 0.73 0.74 0.68 0.70
Llama70b 0.77 0.71 0.86 0.78
Llama70b persona 0.77 0.72 0.83 0.77
Llama70b w target 0.77 0.72 0.86 0.78
D. Llama70b 0.75 0.69 0.84 0.76
D. Llama70b persona 0.75 0.69 0.87 0.77
D. Llama70b w target 0.76 0.71 0.83 0.76
Llama8b 0.73 0.66 0.88 0.76
Llama8b persona 0.74 0.68 0.88 0.76
Llama8b w target 0.74 0.67 0.89 0.76
D. Llama8b 0.70 0.63 0.90 0.74
D. Llama8b persona 0.68 0.60 0.92 0.73
D. Llama8b w target 0.71 0.64 0.90 0.75
o4-mini 0.78 0.71 0.90 0.79
o4-mini persona 0.78 0.70 0.91 0.79
o4-mini w target 0.78 0.71 0.92 0.80
Table 7: The effect of target group on toxicity detection
Appendix J Accuracy Per Higher Target Group
Figure 5: Accuracy For each Higher Target Group
Appendix K Target Group Accuracy Per Target Group
Figure 6: Target group prediction accuracy for each target group